im trying to find where the center of mass of a rectangular table made up from a thin homogeneous plate of 1.2m by 2.3m weighing 8.5kg. It has 4 thin rod shaped legs on its corners. They are homogeneous and have a height of 58cm and weigh 1,7kg each.
1
Expert's answer
2019-11-11T15:51:31-0500

From symmetry, center of mass situated on vertical axis which parallel to rods, and cross center of plate.

Center of mass is beneath plate. Denote mass of plate "M=8.5\text{ kg}" , mass of rod "m=1.7\text{ kg}" and its height "H=0.58 \text{ m}".

Find distance from center of plate to center of mass:


"h = \frac{M\cdot 0 + 4m\cdot H}{M+4m} = \frac{4m\cdot H}{M+4m}=0.26\text{ m} = 36 \text{ cm}"

Finally, the center of mass is 26 cm under of the center of plate.